A light plane must reach a speed of 40m/s for takeoff.

Part A

How long a runway is needed if the (constant) acceleration is 2.8m/s2 ?
Answer #1

v^2 = v0^2 + 2 a x

40^2 = 0^2 + 2 *2.8*x

x = 40^2/(2*2.8)= 285.7 m
Answer #2

We know,

v^{2}=u^{2}+2ad

here u=0 as planes usually starts from rest for take off.

40^{2}=0^{2}+2\times 2.8d

d=285.71 m